OpenStudy (anonymous):

if angle ABC =125 and angle 2 = 40 then find angle 1

OpenStudy (anonymous):

all angles in a triangle add up to 180. If two of your angles are 125 and 40, then your last should be 15.

OpenStudy (anonymous):

Let us consider this as a triangle. The sum of the measures of the angles of a triangle is 180, thus,125+40+x=180 x there is the angle 1. 165+x=180 x=180-165 x=15

OpenStudy (anonymous):

my options are 55 95 85 or 165

OpenStudy (anonymous):

maybe I misunderstood the problem. Does it say that one angle is 125, and another angle is 40?

OpenStudy (anonymous):

can you illustrate the problem please?the information is lacking actually.

OpenStudy (anonymous):

its a obtuse angle that splits in 2 angle 2 = 40 (acute) and ABC as a whole is 125 i'm missing angle 1

OpenStudy (anonymous):

all you have to do is let x be the angle 1. and, x+40=125 x=125-40 x=85

OpenStudy (anonymous):

angle 1 is 85 degrees.